OpenStudy (anonymous):

Which is another way to write the expression 12 • (r – 6)? A. (12 • r) – (12 • 6) B. 12 • r – 6 C. (12 • r) • 6 D. (12 • r) – 6

OpenStudy (anonymous):

I need despert help

OpenStudy (anonymous):

@mossyfish

OpenStudy (anonymous):

I just need a simple answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let's expand this. We have 12*r and 12*-6. Which answer matches this?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@mossyfish

OpenStudy (anonymous):

Excellent!! :)

OpenStudy (anonymous):

What value for b makes this equation true? 2 • 37 = (2 • b) + (2 • 7)

OpenStudy (anonymous):

What about this @mossyfish

OpenStudy (anonymous):

Let's solve this as far as we can, first. Starting with the left side, what is 2*37?

OpenStudy (anonymous):

54

OpenStudy (anonymous):

No. 2*30=60, right? 2*7=14, right? 14+60=74. 2*37=74. This means that, to balance the equation, everything on the right side must equal 74, right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Alright, so now let's solve the right side :). We can't solve the first part because we don't know what b is. Let's look at the second part. 2*7. This is 14, right? So, what plus 14 is equal to 74?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Right, so what times 2 is equal to 60?

OpenStudy (anonymous):

@Julianm we already got that can u help on the other on here

OpenStudy (anonymous):

Hurry up and reply

OpenStudy (anonymous):

Please don't be rude, I asked you a question. What times 2 is equal to 60?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

Right. So, let's plug this back into the equation to check our work. 2*37=(2*(30))+(2*7). 2*30=60, 2*7=14. 60+14=74. 2*37=74. It works! b=30 :)

OpenStudy (anonymous):

What value for z makes this equation true? z • 5 = (50 • 5) + (6 • 5)

OpenStudy (anonymous):

Try this one on your own, I'll check your work

OpenStudy (anonymous):

Ethan, have you solved it?